# Paired Parton Trial States for the Superfluid-Fractional Chern Insulator Transition

### v1

@@ Abstract
Much of our understanding of fractional quantum Hall (FQH) physics comes from wavefunctions which have very high overlap with exact diagonalization results. For fractional Chern insulators (FCIs) [1, 2, 3], the lattice analog of FQH states, similarly good trial wavefunctions are only known to exist in a few special cases that can be mapped to a FQH problem — in particular for so-called ideal [4] or vortexable [5, 6] bands with zero band width [7]. In this paper, we investigate trial wavefunctions for FCIs suggested by the parton construction for the bosonic $\nu=\frac{1}{2}$ state, demonstrating that they can have very large overlap with the exact ground state — even when the underlying band has significant band width. We show the value of these wavefunctions by using them to examine a particular phase transition in detail.
@@
where we assumed for simpliciy that the two partons are independent at the MF level. This assumption is relaxed in Eq. 5. To understand Eq. 3, which is a free Hamiltonian, we need to address a subtlety relating to translational symmetry. We impose that $H^{\text{B}}$ is invariant under two bosonic translation operators $T_{1,2}^{\text{B}}$ which correspond to translations by the primitive lattice vectors and obey $T^{\text{B}}_{1}T^{\text{B}}_{2}=T^{\text{B}}_{2}T^{\text{B}}_{1}$. We may similarly define translational operators in the parton language $T_{1,2}^{\text{P}}$, but the partons only need to fall into a projective representation of this symmetry, in particular we may have an ansatz where $T^{\text{P}}_{1}T^{\text{P}}_{2}=-T^{\text{P}}_{2}T^{\text{P}}_{1}$ [21, 23]. A physical intuition for this comes from Ref. [11] where a Landau level is considered instead of a lattice $\mathcal{C}=1$ band. In that case, the unit-charge boson sees $2\pi$ flux per plaquette, but Eq. 2 dictates that the two partons share the boson’s charge. Assuming they each have half-unit charge, they each see $\pi$-flux, leading to the above magnetic translation relations for $T^{\text{P}}_{1,2}$, effectively doubling the unit-cell. The fact that both partons see equal charge is a non-trivial assumption about the nature of the parton MF state. The success of our approach (below) confirms such an assumption is sensible, but we should note that it is far from guaranteed — for example for the same Hamiltonian, but for filling $\nu=\frac{2}{3}$, the two partons would see different charges and we would get a tripling of the unit cell instead. For $\nu=\frac{1}{2}$ we checked that other extensions of the unit cell do not lead to an improvement in the variational wavefunction. The magnetic translation relations for $T^{\text{P}}_{1,2}$ mean that the partons see a doubled unit cell which we take to be generated by the commuting $(T^{\text{P}}_{1})^{2}$ and $T^{\text{P}}_{2}$ [11, 23]. Since the initial problem was at half-filling, each flavor of partons fills a band at the MF level. The Chern numbers of these bands determine the phase of the bosonic system. If $t(\mathbf{x},\mathbf{y})$ and the expectation value in Eq. 3 are such that the partons have Chern numbers $\mathcal{C}^{\text{P}}_{1}=\mathcal{C}^{\text{P}}_{2}=1$, it can be shown that the system is in a FCI phase. The phase $\mathcal{C}^{\text{P}}_{1}=1$, $\mathcal{C}^{\text{P}}_{2}=0$ is a Mott insulator and $\mathcal{C}^{\text{P}}_{1}=1$, $\mathcal{C}^{\text{P}}_{2}=-1$ is a superfluid [11]. While the argument in Ref. [11] is field-theoretical, much of this may be understood in terms of wavefunctions.
@@
We might first try to optimize over all possible ansatze of the determinant-product form described below Eq. 4 as this general form can capture both sides of the transition. While such a procedure sees a transition, we find it to be a relatively poor description of the superfluid phase when comparing to exact diagonalization, as we show in Fig. 1. The reason for this may be traced back to the fact that the partons are assumed independent in the MF state, so $\left\langle b\right\rangle_{\text{MF}}=\left\langle f_{1}f_{2}\right\rangle_{\text{MF}}=0$ in that state. But this goes against the usual idea that this symmetry-broken phase should be understood in terms of fluctuations $\delta b=b-\langle b\rangle$ around $\langle b\rangle\neq 0$. To obtain good results for the transition, we must first find an accurate description of the superfluid state. Based on the argument above, we should put the partons in a paired mean-field state which should at least allow for a nonzero $\left\langle b\right\rangle_{\text{MF}}=\left\langle f_{1}f_{2}\right\rangle_{\text{MF}}$.
We note here the distinction between the MF and true expectation values — for an operator $\mathcal{O}$ and a MF state $\left.\middle|\Omega\right\rangle$, we have $\left\langle\mathcal{O}\right\rangle_{\text{MF}}=\left\langle\Omega\middle|\mathcal{O}\middle|\Omega\right\rangle$ while the true expectation value in the state $\left.\middle|\phi\right\rangle\propto P\left.\middle|\Omega\right\rangle$ projected onto the physical subspace (eg. Eq. 4) is different. In operator language, $P=\prod_{\mathbf{r}}\left[\hat{n}_{1}(\mathbf{r})\hat{n}_{2}(\mathbf{r})+(1-\hat{n}_{1}(\mathbf{r}))(1-\hat{n}_{2}(\mathbf{r}))\right]$. Any valid bosonic operator (incl. $H^{\text{B}}$) must commute with the projector $P$  [^1] This translates to requiring $\mathcal{O}$ to be a singled w.r.t. the $SU(2)$ gauge field in the field-theoretical approach.. This, together with $P^{2}=P$, implies for the projected expectations $\left\langle\phi\middle|\mathcal{O}\middle|\phi\right\rangle=\langle\mathcal{O}\rangle=\left\langle\Omega\middle|\mathcal{O}P\middle|\Omega\right\rangle/\left\langle\Omega\middle|P\middle|\Omega\right\rangle=\frac{\left\langle\mathcal{O}P\right\rangle_{\text{MF}}}{\left\langle P\right\rangle_{\text{MF}}}$. Thus the determinant product state below Eq. 4 can have $\left\langle b\right\rangle\neq 0$ after projection despite having $\left\langle b\right\rangle_{\text{MF}}=0$. The MF expectations evaluate operators on both the physical and unphysical subspaces — still, certain MF quantities such as Chern numbers may be useful to interpret the projected state.
@@
for some $t^{\text{MF}},\Delta^{\text{MF}}$. Considering possible parton contractions in Eq. 1, it may seem that many of the terms in Eq. 5 must be zero. But the operator we should really consider is not $H^{\text{B}}$, but $H^{\text{B}}P$. This will lead to more contractions, possibly over larger distances. Instead of explicitly evaluating the coefficients in Eq. 5, we note that the ground state will always be a Pfaffian. Say we have $M$ partons ($M=2N$ for $N$ bosons) at positions $\mathbf{r}_{j}$ and of flavors $\alpha_{j}$ for $j=1\ldots M$. Then, the ground state is given by the Pfaffian form [26]
@@
for some function $g_{\alpha\beta}(\mathbf{x},\mathbf{y})=-g_{\beta\alpha}(\mathbf{y},\mathbf{x})$. Due to the possible presence of terms such as $f_{1}^{\dagger}f_{2}$ in Eq. 5, the partons may change flavor. We can think of flavor as an additional site index — each physical site $\mathbf{r}$ splits into $K$ sub-sites indexed by $(\mathbf{r},\alpha)$ for $\alpha=1\ldots K$. The parton ‘flavor’ then simply specifies at which of these sub-sites the parton resides. In this language, the physical subspace corresponds to states where each physical site either has all $K$ of its sub-sites occupied or all of them empty. In our $K=2$ case, a state with $N$ bosons at positions $\mathbf{r}_{1},\ldots,\mathbf{r}_{N}$ is described by filling both sub-sites at each of those sites in Eq. 6, giving the bosonic wavefunction
@@
This obeys both bosonic statistics and the hard-core constraint. It is equivalent to using a paired MF state $|\tilde{\Omega}\rangle$ in Eq. 4. Our method in what follows is to take the form of Eq. 7 as our wavefunction ansatz, using the values of the pairing function $g_{\alpha\beta}(\mathbf{x},\mathbf{y})$ as our variational parameters. As $\phi$ is an ansatz for the full bosonic wavefunction, the optimization scheme will be to directly minimize $E[\phi]=\frac{\left\langle\phi\middle|H^{\text{B}}\middle|\phi\right\rangle}{\left\langle\phi\middle|\phi\right\rangle}$, bypassing the need to evaluate $H^{\text{MF}}$ in Eq. 5. The evaluation of $E[\phi]$ is however not analytically tractable, forcing us into variational Monte-Carlo (VMC) sampling. Details of the sampling and optimization are shown in the Supplement [20], Sect. II. Our VMC procedure is different from any mean-field scheme (e.g. solving Eq. 3 self-consistently) as it optimizes $\langle H^{\text{B}}\rangle=\left\langle H^{\text{B}}P\right\rangle_{\text{MF}}/\left\langle P\right\rangle_{\text{MF}}$ and not $\left\langle H^{\text{B}}\right\rangle_{\text{MF}}$.
We require that the function $g_{\alpha\beta}(\mathbf{x},\mathbf{y})$ in Eq. 6 obeys the translational symmetry mandated by the larger, partonic unit cell (so $(T^{\text{P}}_{1})^{2}$ and $T^{\text{P}}_{2}$). We do not enforce any action of $T^{\text{P}}_{1}$ itself on $g_{\alpha\beta}(\mathbf{x},\mathbf{y})$ — this allows the total bosonic-unit-cell momentum to be $\mathbf{Q}=(0,0)$ and $\mathbf{Q}=(\pi,0)$. By considering $\phi(\{\mathbf{r}\})\pm\phi(\{\mathbf{\tilde{r}}\})$ with $\{\mathbf{\tilde{r}}\}$ having all particles translated by $T_{1}^{\text{B}}$ relative to $\{\mathbf{r}\}$, we can construct an ansatz for each of the above $\mathbf{Q}$. For system dimensions where the two topologically degenerate states in the FCI phase occur at different momenta, this will allow us to generate both states.
@@
Computing overlaps, our results are remarkable — in a Hilbert space of dimension $10^{8}$, so $\dim\mathcal{H}\sim 5\cdot 10^{6}$ per momentum, we find a $99\%$ overlap in the flatband limit and an overlap $\gtrsim 91\%$ throughout the entire FCI-SF transition for the $\mathbf{Q}=(0,0)$ VMC ansatz. The function $g_{\alpha\beta}(\mathbf{x},\mathbf{y})$ over which we optimize contains $d=567$ independent parameters. This gives $\frac{\dim\mathcal{H}}{d}\sim 10^{4}$ and the high overlaps indicate that the WF ansatz form captures the physics well. We show energies for both the best paired VMC state at $\mathbf{Q}=(0,0)$ and at $\mathbf{Q}=(\pi,0)$. While the latter is equally good in the FCI limit, it is high in energy in the SF phase because of its momentum. We also show the energies of the determinant-product state — that is assuming independent partons but still using the full VMC technology. While the energy is good in the flatband limit and the wavefunction only has $d^{\prime}=54$ parameters, Fig. 1 shows that this ansatz fails to captures the energetics of the SF state, demonstrating the need for the paired ansatz. In Sect. III of the Supplement [20] we also compare our method to the DMRG results of Ref. [10] on the honeycomb lattice, with similar conclusions.
@@
With a periodic structure in $g_{\alpha\beta}(\mathbf{x},\mathbf{y})$, we may define a MF BdG parton Chern number $\mathcal{C}^{\text{BdG}}$ as discussed in the Supplement [20], Sect. IV. In the limit of decoupled partons, we have $\mathcal{C}^{\text{BdG}}=2(\mathcal{C}^{\text{P}}_{1}+\mathcal{C}^{\text{P}}_{2})$, suggesting $\mathcal{C}^{\text{BdG}}=4$ for FCI and $\mathcal{C}^{\text{BdG}}=0$ for SF. A jump of $\mathcal{C}^{\text{BdG}}$ by four units is natural — by the BdG nature, a gap closure at $\mathbf{k}$ implies one at $-\mathbf{k}$ and the projective translation argument then necessitates ±$\pm\mathbf{k}+(0,\pi)$, generically giving four closures as in Ref. [13]. We observe such a jump at the transition, directly confirming the theoretical proposal of Refgs. [11, 14] with the transition described by a multiple parton gap closure. But our numerics show that to accurately capture the physics, a nonzero $\left\langle f_{1}f_{2}\right\rangle_{\text{MF}}$ must be allowed. It remains to be seen how this affects the field theory of the transition from [11]. In Supplement[20], Sect. V, we present a self-consistent MF method, informed by our above results which allows us to work with systems large enough to clearly see the quadruple band closure. There are scenarios (explained in detail in Supplement [20], Sect. IV) where $\mathcal{C}^{\text{BdG}}$ can change in steps of 2 without breaking translational symmetry and thus we can have a phase with $\mathcal{C}^{\text{BdG}}=2$ which is different from what has been proposed previously, as it can only emerge from a paired-parton ansatz. However, our numerics have not found examples where this is energetically favorable.
We have so far considered models with the on-site hard-core constraint being the only interaction. We now add a repulsive density-density interaction $\hat{V}=\frac{1}{2}V_{1}\sum_{\langle\mathbf{x},\mathbf{y}\rangle}\hat{n}_{\text{B}}(\mathbf{x})\hat{n}_{\text{B}}(\mathbf{y})$ between neighboring sites $\langle\mathbf{x},\mathbf{y}\rangle$. We show the energy and WF overlap agreement with exact results in Fig. 2 as a function of the interaction strength $V_{1}$.
@@
While the agreement is good for relatively weak $V_{1}\lesssim\text{max}(|t|)=1$, at $V_{1}\sim 1$ our trial states have an energy above the gap and a relatively poor ground-state overlap. Throughout the variation, the two topological ground states remain degenerate and our VMC procedure generates both equally well. Finally, we note that for large $V_{1}$, this system approaches a supersolid phase [17, 27, 28], which may be part of the reason why our ansatz works relatively poorly there.

### v2

@@ Abstract
Much of our understanding of fractional quantum Hall (FQH) physics comes from wavefunctions which have very high overlap with exact diagonalization results. For fractional Chern insulators (FCIs) [21, 4, 13, 22, 36, 16], the lattice analog of FQH states, similarly good trial wavefunctions are only known to exist in a few special cases that can be mapped to a FQH problem — in particular for so-called ideal [25] or vortexable [11, 7] bands with zero band width [9]. In this paper, we investigate trial wavefunctions for FCIs suggested by the parton construction for the bosonic $\nu=\frac{1}{2}$ state, demonstrating that they can have very large overlap with the exact ground state — even when the underlying band has significant band width. We show the value of these wavefunctions by using them to examine a particular phase transition in detail.
@@
where we assumed for simpliciy that the two partons are independent at the MF level. This assumption is relaxed in Eq. 5. To understand Eq. 3, which is a free Hamiltonian, we need to address a subtlety relating to translational symmetry. We impose that $H^{\text{B}}$ is invariant under two bosonic translation operators $T_{1,2}^{\text{B}}$ which correspond to translations by the primitive lattice vectors and obey $T^{\text{B}}_{1}T^{\text{B}}_{2}=T^{\text{B}}_{2}T^{\text{B}}_{1}$. We may similarly define translational operators in the parton language $T_{1,2}^{\text{P}}$, but the partons only need to fall into a projective representation of this symmetry, in particular we may have an ansatz where $T^{\text{P}}_{1}T^{\text{P}}_{2}=-T^{\text{P}}_{2}T^{\text{P}}_{1}$ [35, 18]. A physical intuition for this comes from Ref. [3] where a Landau level is considered instead of a lattice $\mathcal{C}=1$ band. In that case, the unit-charge boson sees $2\pi$ flux per plaquette, but Eq. 2 dictates that the two partons share the boson's charge. Assuming they each have half-unit charge, they each see $\pi$-flux, leading to the above magnetic translation relations for $T^{\text{P}}_{1,2}$, effectively doubling the unit-cell. The fact that both partons see equal charge is a non-trivial assumption about the nature of the parton MF state. The success of our approach (below) confirms such an assumption is sensible, but we should note that it is far from guaranteed — for example for the same Hamiltonian, but for filling $\nu=\frac{2}{3}$, the two partons would see different charges and we would get a tripling of the unit cell instead. For $\nu=\frac{1}{2}$ we checked that other extensions of the unit cell do not lead to an improvement in the variational wavefunction. The magnetic translation relations for $T^{\text{P}}_{1,2}$ mean that the partons see a doubled unit cell which we take to be generated by the commuting $(T^{\text{P}}_{1})^{2}$ and $T^{\text{P}}_{2}$ [3, 18]. Since the initial problem was at half-filling, each flavor of partons fills a band at the MF level. The Chern numbers of these bands determine the phase of the bosonic system. If $t(\mathbf{x},\mathbf{y})$ and the expectation value in Eq. 3 are such that the partons have Chern numbers $\mathcal{C}^{\text{P}}_{1}=\mathcal{C}^{\text{P}}_{2}=1$, it can be shown that the system is in a FCI phase. The phase $\mathcal{C}^{\text{P}}_{1}=1$, $\mathcal{C}^{\text{P}}_{2}=0$ is a Mott insulator and $\mathcal{C}^{\text{P}}_{1}=1$, $\mathcal{C}^{\text{P}}_{2}=-1$ is a superfluid [3]. While the argument in Ref. [3] is field-theoretical, much of this may be understood in terms of wavefunctions.
@@
We might first try to optimize over all possible ansatze of the determinant-product form described below Eq. 4 as this general form can capture both sides of the transition. While such a procedure sees a transition, we find it to be a relatively poor description of the superfluid phase when comparing to exact diagonalization, as we show in Fig. 1. The reason for this may be traced back to the fact that the partons are assumed independent in the MF state, so $\left\langle b\right\rangle_{\text{MF}}=\left\langle f_{1}f_{2}\right\rangle_{\text{MF}}=0$ in that state. Such MF-independent partons can display a superfluid phase [3], but the resulting SF wavefunction is not accurate. In Sect. III of [26], we develop a field-theoretical picture showing that parton states with a nonzero $\langle f_{1}f_{2}\rangle_{\text{MF}}$ describe the same SF phase and that the MF saddle-point generally has a nonzero $\langle f_{1}f_{2}\rangle_{\text{MF}}$. This motivates considering paired parton trial states.
We note here the distinction between the MF and true expectation values — for an operator $\mathcal{O}$ and a MF state $\left.\middle|\Omega\right\rangle$, we have $\left\langle\mathcal{O}\right\rangle_{\text{MF}}=\left\langle\Omega\middle|\mathcal{O}\middle|\Omega\right\rangle$ while the true expectation value in the state $\left.\middle|\phi\right\rangle\propto P\left.\middle|\Omega\right\rangle$ projected onto the physical subspace (eg. Eq. 4) is different. In operator language, $P=\prod_{\mathbf{r}}\left[\hat{n}_{1}(\mathbf{r})\hat{n}_{2}(\mathbf{r})+(1-\hat{n}_{1}(\mathbf{r}))(1-\hat{n}_{2}(\mathbf{r}))\right]$. Any valid bosonic operator (including $H^{\text{B}}$) must commute with the projector $P$  [^1] This translates to requiring $\mathcal{O}$ to be a singlet w.r.t. the $SU(2)$ gauge field in the field-theoretical approach.. This, together with $P^{2}=P$, implies for the projected expectations $\left\langle\phi\middle|\mathcal{O}\middle|\phi\right\rangle=\langle\mathcal{O}\rangle=\left\langle\Omega\middle|\mathcal{O}P\middle|\Omega\right\rangle/\left\langle\Omega\middle|P\middle|\Omega\right\rangle=\frac{\left\langle\mathcal{O}P\right\rangle_{\text{MF}}}{\left\langle P\right\rangle_{\text{MF}}}$. The MF expectations evaluate operators on both the physical and unphysical subspaces — still, MF quantities such as Chern numbers are useful to interpret the projected state.
@@
for some $t^{\text{MF}},\Delta^{\text{MF}}$. Considering possible parton contractions in Eq.§ 1, it may seem that many of the terms in Eq. 5 must be zero. But the operator we should really consider is not $H^{\text{B}}$, but $H^{\text{B}}P$. This will lead to more contractions, possibly over larger distances. While this allows for any fermionic bilinear to have a nonzero MF expectation, depending on $H^{\text{B}}$, the best variational wavefunctions may occur when some bilinears have $\langle\cdot\rangle_{\text{MF}}=0$. Instead of explicitly evaluating the coefficients in Eq. 5, we note that the ground state, projected to a fixed number of bosons, will always be a Pfaffian. Say we have $M$ partons ($M=2N$ for $N$ bosons) at positions $\mathbf{r}_{j}$ and of flavors $\alpha_{j}$ for $j=1\ldots M$. Then, the ground state is given by the Pfaffian form [24]
@@
for some function $g_{\alpha\beta}(\mathbf{x},\mathbf{y})=-g_{\beta\alpha}(\mathbf{y},\mathbf{x})$. Due to the possible presence of terms such as $f_{1}^{\dagger}f_{2}$ in Eq. 5, the partons may change flavor. We can think of flavor as an additional site index — each physical site $\mathbf{r}$ splits into $K$ sub-sites indexed by $(\mathbf{r},\alpha)$ for $\alpha=1\ldots K$. The parton `flavor' then simply specifies at which of these sub-sites the parton resides. In this language, the physical subspace corresponds to states where each physical site either has all $K$ of its sub-sites occupied or all of them empty. In our $K=2$ case, a state with $N$ bosons at positions $\mathbf{r}_{1},\ldots,\mathbf{r}_{N}$ is described by filling both sub-sites at each of those sites in Eq. 6, giving the bosonic wavefunction
@@
This obeys both bosonic statistics and the hard-core constraint. It is equivalent to using a paired MF state $|\tilde{\Omega}\rangle$ in Eq. 4. Our method in what follows is to take the form of Eq. 7 as our wavefunction ansatz, using the values of the pairing function $g_{\alpha\beta}(\mathbf{x},\mathbf{y})$ as our variational parameters. As $\phi$ is an ansatz for the full bosonic wavefunction, the optimization scheme will be to directly minimize $E[\phi]=\frac{\left\langle\phi\middle|H^{\text{B}}\middle|\phi\right\rangle}{\left\langle\phi\middle|\phi\right\rangle}$, bypassing the need to evaluate $H^{\text{MF}}$ in Eq. 5. The evaluation of $E[\phi]$ is however not analytically tractable, forcing us into variational Monte-Carlo (VMC) sampling. Details of the sampling and optimization are shown in the Supplement [26], Sect. IV. Our VMC procedure is different from any mean-field scheme (e.g. solving Eq. 3 self-consistently) as it optimizes $\langle H^{\text{B}}\rangle=\left\langle H^{\text{B}}P\right\rangle_{\text{MF}}/\left\langle P\right\rangle_{\text{MF}}$ and not $\left\langle H^{\text{B}}\right\rangle_{\text{MF}}$.
[…]
@@
Computing overlaps, our results are remarkable — in a Hilbert space of dimension $10^{8}$, so $\dim\mathcal{H}\sim 5\cdot 10^{6}$ per momentum, we find a $99\%$ overlap in the flatband limit and an overlap $\gtrsim 91\%$ throughout the entire FCI-SF transition for the $\mathbf{Q}=(0,0)$ VMC ansatz. The function $g_{\alpha\beta}(\mathbf{x},\mathbf{y})$ over which we optimize contains $d=567$ independent parameters. This gives $\frac{\dim\mathcal{H}}{d}\sim 10^{4}$ and the high overlaps indicate that the WF ansatz form captures the physics well. We show energies for both the best paired VMC state at $\mathbf{Q}=(0,0)$ and at $\mathbf{Q}=(\pi,0)$. While the latter is equally good in the FCI limit, it is high in energy in the SF phase because of its momentum. We also show the energies of the determinant-product state — that is assuming independent partons but still using the full VMC technology. While the energy is good in the flatband limit and the wavefunction only has $d^{\prime}=54$ parameters, Fig. 1 shows that this ansatz fails to captures the energetics of the SF state, demonstrating the need for the paired ansatz. In Sect. V of the Supplement [26] we also compare our method to the DMRG results of Ref. [15] on the honeycomb lattice, with similar conclusions. While overlaps are considered to be a gold-standard for comparison of wavefunctions to numerics, finite size effects should be treated very carefully. We have seen no indication of behavioral change at larger systems sizes and in Sect. II of the Supplement [26], we show how similar wavefunctions can be generated for arbitrary systems sizes without indication of any new gap closings.
We find numerically that only terms of the form $\langle f_{1}^{\dagger}f_{1}\rangle_{\text{MF}}$$\langle f_{2}^{\dagger}f_{2}\rangle_{\text{MF}}$ and $\langle f_{1}f_{2}\rangle_{\text{MF}}$ obtain a significant nonzero expectation value. So while the ansatz Eq. 5 is able to explore a variety of mean-field states, including ones with the gauge group broken to $\mathbb{Z}_{2}$, our results suggest that states with a $U(1)$ gauge group are favoured in this model. Our FCI state is topologically identical to that of Ref. [3]. As discussed in Sect. III of [26], the nonzero $\langle f_{1}f_{2}\rangle_{\text{MF}}$ keeps us in the same SF phase, but improves energetics and allows for a larger SF weight. In fact in Sect. II of the Supplement [26], we show that most of the results in Fig. 1 can be recovered by an explicitly $U(1)$-invariant short-range parton MF Hamiltonian analogous to Eq. 5.
@@
With a periodic structure in $g_{\alpha\beta}(\mathbf{x},\mathbf{y})$, we may define a MF BdG parton Chern number $\mathcal{C}^{\text{BdG}}$ as discussed in the Supplement [26], Sect. VI. In the limit of decoupled partons, we have $\mathcal{C}^{\text{BdG}}=2(\mathcal{C}^{\text{P}}_{1}+\mathcal{C}^{\text{P}}_{2})$, suggesting $\mathcal{C}^{\text{BdG}}=4$ for FCI and $\mathcal{C}^{\text{BdG}}=0$ for SF. A jump of $\mathcal{C}^{\text{BdG}}$ by four units is natural — by the BdG nature, a gap closure at $\mathbf{k}$ implies one at $\mathbf{k}$ and the projective translation argument then necessitates $\pm\mathbf{k}+(0,\pi)$, generically giving four closures as in Ref. [2]. We observe such a jump at the transition, directly confirming the theoretical proposal of Refs. [3, 28] with the transition described by a multiple parton gap closure. But our numerics show that to accurately capture the physics, a nonzero anomalous expectation $\Delta=\left\langle f_{1}f_{2}\right\rangle_{\text{MF}}$ must be allowed. While $\Delta$ is expected to be zero at the transition in the thermodynamic limit, we argue in Sect. III of [26] that its fluctuations affect details of the transition, including a re-normalization of the transition point, which we also observe numerically in an extrapolation to large system sizes.
In Supplement[26], Sect. VII, we present a self-consistent MF method, informed by our above results which allows us to work with systems large enough to clearly see the quadruple band closure. There are scenarios (explained in detail in Supplement [26], Sect. VI) where $\mathcal{C}^{\text{BdG}}$ can change in steps of 2 without breaking translational symmetry and thus we can have a phase with $\mathcal{C}^{\text{BdG}}=2$ which is different from what has been proposed previously, as it can only emerge from a paired-parton ansatz. However, our numerics have not found examples where this is energetically favorable.
@@
We have so far considered models with the on-site hard-core constraint being the only interaction. We now add a repulsive density-density interaction $\hat{V}=\frac{1}{2}V_{1}\sum_{\langle\mathbf{x},\mathbf{y}\rangle}\hat{n}_{\text{B}}(\mathbf{x})\hat{n}_{\text{B}}(\mathbf{y})$ between neighboring sites $\langle\mathbf{x},\mathbf{y}\rangle$. We show the energy and WF overlap agreement with exact results in Fig. 2 as a function of the interaction strength $V_{1}$. While the agreement is good for relatively weak $V_{1}\lesssim\text{max}(|t|)=1$, at $V_{1}\sim 1$ our trial states have an energy above the gap and a relatively poor ground-state overlap. Throughout the variation, the two topological ground states remain degenerate and our VMC procedure generates both equally well. Finally, we note that for large $V_{1}$, this system approaches a supersolid phase [33, 17, 14], which may be part of the reason why our ansatz works relatively poorly there.
@@
Note added — Since our initial posting of this work, Ref. [32] has appeared online studying the same transition. Our results are consistent where they overlap.